find the equation of the line between (1,4)and(3,1), don't actually need to graph this, i thought this would be the best section for the question.
Thanks in advance.

Answer by mananth(16949)

x1 y1 x2 y2
1 4 3 1

slope m =(y2-y1)/(x2-x1)
(1-4)/(3-1)
(-3/2)
m=-1.5

Plug value of the slope and point (1, 4) in
Y = m x + b
4.00 = -1.5 + b
b= 4 -1.5
b= 5.5
So the equation will be
Y =-1.5x+5.5
m.ananth@hotmail.ca